OpenStudy (anonymous):

What is the area of an equilateral triangle with side lengths 17 cm and an altitude of 14.72 cm?

jimthompson5910 (jim_thompson5910):

use the formula A = (b*h)/2 in this case, the base is 17, so b = 17 also in this case, the height is 14.72, which means h = 14.72

OpenStudy (nurali):

a=17 and b=14.72 area = ab/2 = 17*14.7224/2 --> 125.14

OpenStudy (anonymous):

oh its 125.14? ..the answer I came up with was 125.12 ;)

jimthompson5910 (jim_thompson5910):

A = (b*h)/2 A = (17*14.72)/2 A = (250.24)/2 A = 125.12 So if you use the altitude of 14.72, then you will get an answer of 125.12

jimthompson5910 (jim_thompson5910):

it turns out that the altitude is more like (17/2)*sqrt(3) = 14.7224318643354 (roughly), so that explains how Nurali got the number 14.7224

OpenStudy (anonymous):

alright thank you its actually pretty simple now that you guys explained it

jimthompson5910 (jim_thompson5910):

you're welcome